OpenStudy (anonymous):

the cost of a small yogurt is $2.50 plus $0.15 per topping. The equation that gives you the total cost c of a small frozen yogurt is c=0.15+2.50, where n is the number of toppings. What is the cost of a small frozen yogurt with 3 toppings?

OpenStudy (madhu.mukherjee.946):

c=2.50+0.45 where n=3

OpenStudy (madhu.mukherjee.946):

so cost of toppings =n*o.15 =3*0.15 =0.45

OpenStudy (madhu.mukherjee.946):

c=2.95

OpenStudy (anonymous):

Thankyou!

OpenStudy (madhu.mukherjee.946):

welcome